OpenStudy (anonymous):

Can anyone help with this one please N+ 8.7 = 16.2

OpenStudy (ahvei):

N=7.5

OpenStudy (ahvei):

medal plsss

OpenStudy (anonymous):

What does that mean

OpenStudy (anonymous):

Im new to this

OpenStudy (ahvei):

nvm then

OpenStudy (anonymous):

The sum of constant N and 8.7 equals to 16.2--- N+8.7=16.2 Therefore, N=16.2-8.7 N=7.5